Janet invested $27,000, part at 15% and part at 9%. If the total interest at the end of the year is $3,390, how much did she invest at 15%?
Answer by stanbon(57224)

Janet invested $27,000, part at 15% and part at 9%. If the total interest at the end of the year is $3,390, how much did she invest at 15%?
---------------------
Using two variables:
Equations:
Quantity Equation: x + y = 27000
Interest Equation:0.15x + 0.09y = 3390
----------------------------------------
Multiply thru the 1st equation by 9
Multiply thru the 2nd equation by 100
-----
9x + 9y = 9*27000
15x+ 9y = 339000
------------------------
Subtract the 1st from the 2nd and solve for "x":
6x = 96000
x = $16,000 (amt. invested at 15 %
------------------------------
Cheers,
Stan H.